OpenStudy (anonymous):

what is x^2 over x^2

OpenStudy (solomonzelman):

\(\Huge\color{blue}{ \bf \frac{x^{2}}{x^{2}}=1 }\)

OpenStudy (anonymous):

So that would be 1 and not just x?

OpenStudy (solomonzelman):

When you have a NUMBER/NUMBER it will always equal 1 (besides 0)

OpenStudy (solomonzelman):

No, not just x. \(\Huge\color{blue}{ \bf \frac{x^2}{x}=0 }\)

OpenStudy (solomonzelman):

but x^2/x^2 =1

OpenStudy (anonymous):

1 is ans

OpenStudy (primeralph):

@SolomonZelman Why 0?

OpenStudy (solomonzelman):

I meant to say x... I am sorry :)